Prove that 1 horsepower= 746 watts using concepts of current electricity.

Grade:12

2 Answers

Vikas TU
7 years ago
one foot =0.3408 metres.
 
1 HP = 550 Foot Pound per second.
1 Pound = 4.44822 meters.
1 Foot is 0.3408 meters.
1 Hp = 550*4.44822*0.3048 meters
     = 745.69 Nm/s = 746 Watts.
Soham
7 years ago
1HP = 550 Foot Pound per second 1 Pound = 4.44822 Newtons1 Foot is 0.3048 Meters.1 HP = 550 * 4.44822 * 0.3048 meters1Hp = 745.69 Newton meter per second. A 1 hp motor cannot use less than 745.7 Watts so if it uses 1 kW it is about 75% efficient, which sounds about right for a motor of that size.